OpenStudy (anonymous):

what is the x-intercept of the tangent to y=e^x at x=45

OpenStudy (anonymous):

line tangent to the graph has slope \[e^{45}\]

OpenStudy (anonymous):

and goes through the point \[(45,e^{45})\] so equation is \[y-e^{45}=e^{45}(x-45)\]

OpenStudy (anonymous):

x intercept means set y = 0 and solve \[0-e^{45}=e^{45}(x-45)\] \[-1=x-45 \] \[x=44\]

OpenStudy (anonymous):

check my algebra because i am tired

myininaya (myininaya):

i got x=44 as well

OpenStudy (anonymous):

thank you, because of yur help i just might graudaute